Solve the equation by graphing:
y = x - 2
x + y = 2
x = and y =

Answer by ewatrrr(24785)

 
Hi,
Solve the equation by graphing:
y = x - 2 Green
x + y = 2 Or y = -x + 2 Blue
Ordered pair (x,y): (2,0) the solution for this system of EQs